### v1

# Topological consequences of null-geodesic refocusing and applications to $Z^{x}$ manifolds

Abstract
Let $(M,h)$ be a connected, complete Riemannian manifold, let $x\in M$ and $l>0$. Then $M$ is called a $Z^{x}$ manifold if all geodesics starting at $x$ return to $x$ and it is called a $Y^{x}_{l}$ manifold if every unit-speed geodesic starting at $x$ returns to $x$ at time $l$. It is unknown whether there are $Z^{x}$ manifolds that are not $Y^{x}_{l}$-manifolds for some $l>0$. By the Bérard-Bergery theorem, any $Y^{x}_{l}$ manifold of dimension at least $2$ is compact with finite fundamental group. We prove the same result for $Z^{x}$ manifolds $M$ for which all unit-speed geodesics starting at $x$ return to $x$ in uniformly bounded time. We also prove that any $Z^{x}$ manifold $(M,h)$ with $h$ analytic is a $Y^{x}_{l}$ manifold for some $l>0$. We start by defining a class of globally hyperbolic spacetimes (called observer-refocusing) such that any $Z^{x}$ manifold is the Cauchy surface of some observer-refocusing spacetime. We then prove that under suitable conditions the Cauchy surfaces of observer-refocusing spacetimes are compact with finite fundamental group and show that analytic observer-refocusing spacetimes of dimension at least $3$ are strongly refocusing. We end by stating a contact-theoretic conjecture analogous to our results in Riemannian and Lorentzian geometry.
journal: Journal of Geometry and Physics
[…]
[inst1]organization=Dartmouth College Department of Mathematics
1 Introduction
Throughout the paper, all Riemannian manifolds and Lorentzian manifolds will be assumed to be smooth, connected, and without boundary. Further all geodesics will be assumed to be maximal unless otherwise stated. If $(X,g)$ is a Lorentzian manifold, $p\in X$ and $v\in T_{p}X$, then $\alpha_{v}$ will always denote the geodesic defined through $\alpha(0)=p$ and $\alpha^{\prime}(0)=v$.
1.1 Definition ($Y^{(x,y)}_{l}$ manifold).
Let $(M,h)$ be a complete Riemannian manifold, $x\in M$ and $l>0$. We say that $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold if for all geodesics $\alpha$ in $(M,h)$ with $\alpha(0)=x$ and $g(\alpha^{\prime}(0),\alpha^{\prime}(0))=1$ we have $\alpha(l)=y$. If $x=y$ we call $(M,h)$ a $Y^{x}_{l}$ manifold. Note also that every $Y^{(x,y)}_{l}$ is a $Y^{x}_{2l}$ manifold (see [12]).
1.2 Definition ($Z^{(x,y)}$ manifold).
Let $(M,h)$ be a complete Riemannian manifold, $x\in M$ We say that $(M,h)$ is a $Z^{(x,y)}$ manifold if for all geodesics $\alpha$ in $(M,h)$ with $\alpha(0)=x$ there is some $t>0$ with $\alpha(t)=y$. If $x=y$ we call $(M,h)$ a $Z^{x}$ manifold. In contrast to $Y^{(x,y)}_{l}$ manifolds it is not known whether every $Z^{(x,y)}$ manifold is a $Z^{x}$ manifold.
1.3 Definition (strongly refocusing spacetime).
Let $(X,g)$ be a spacetime. We say that $(X,g)$ is strongly refocusing if there are $p,q\in X$ with $p\neq q$ such that all null-geodesics $\alpha$ in $(X,g)$ going through $p$ also go through $q$.
1.4 Definition (observer refocusing spacetime).
Let $(X,g)$ be a spacetime, $p\in X$ and $\gamma$ some timelike curve in $X$. We say that $(X,g)$ is observer-refocusing if for all future-oriented null-geodesics $\alpha$ in $(X,g)$ with $\alpha(0)=x$ there is some $t>0$ with $\alpha(t)\in\mathrm{im}(\gamma)$.
Observer-refocusing spacetimes have the following physical interpretation. The worldline of physical observers through spacetime is a timelike curve. A spacetime is observer-refocusing, if there is some point in spacetime and some observer, such that if light is emitted from the point in spacetime in every direction, the observer will, over the course of their life, be able to see all that light. Every $Y^{(x,y)}_{l}$ manifold is a $Z^{(x,y)}$ manifold and every strongly refocusing spacetime is an observer-refocusing spacetime. (It turns out that the roles of $p$ and $q$ in Definition 1.3 are symmetric, see [16], hence we can assume without loss of generality that $q\in J^{+}(p)$ and pick $\gamma$ to be any timelike curve through $q$.)
1.5 Definition (spacetime associated to a Riemannian manifold).
Let $(M,h)$ be a Riemannian manifold. Let $X=M\times\mathbb{R}$ and let $g=h\oplus-dt^{2}$. Then $(X,g)$ is a Lorentzian manifold. $(X,g)$ is a special case of what is more generally known as a warped product of $(M,h)$ and $(\mathbb{R},-dt^{2})$, in this case with trivial warping function. We equip $X$ with the time-orientation given by the vectorfield $\partial_{t}$ to make it a spacetime. We call $(X,g)$ the spacetime associated to the Riemannian manifold $(M,h)$.
1.6 Definition (globally hyperbolic spacetime).
A spacetime $(X,g)$ is called causal if it admits no closed timelike curves. A spacetime is called globally hyperbolic if it is causal and if for all $p,q\in X$ the set $J^{+}(p)\cap J^{-}(q)$ is compact. A submanifold $S\subset X$ is called a Cauchy surface of $(X,g)$ if every inextendible causal curve in $X$ intersects $S$ exactly once. The existence of a Cauchy surface $S$ of a spacetime $(X,g)$ is equivalent to the spacetime being globally hyperbolic, see [15]. Further, any two Cauchy surfaces of a given spacetime $(X,g)$ are diffeomorphic and $X$ is diffeomorphic to $S\times\mathbb{R}$ for any Cauchy surface $S$, see [14].
The existence of Cauchy surfaces ensures that the spacetime is well-behaved and that there are no naked singularities. It turns out that globally hyperbolic spacetimes are exactly the spacetimes obeying the cosmic censorship hypothesis as conjectured by Penrose in [22], making them physically reasonable.
1.7 Definition (temporal functions).
For a spacetime $(X,g)$ a smooth temporal function $\mathcal{T}:X\to\mathbb{R}$ is a smooth function with everywhere future-pointing timelike gradient. In [3] [4] Bernal and Sanchez prove that for every globally hyperbolic spacetime $X$, there are smooth, surjective temporal functions $\mathcal{T}:X\to\mathbb{R}$ such that $S_{t}\coloneqq\mathcal{T}^{-1}(t)\subset X$ is a smooth spacelike Cauchy surface for each $t\in\mathbb{R}$. Smooth temporal functions are strictly increasing along every future-directed causal curve.
1.8 Remark.
If $(M,h)$ is a complete Riemannian manifold, then the associated spacetime $(X,g)$ is globally hyperbolic, geodesically complete, and for each $t\in\mathbb{R}$ the submanifold $M\times\{t\}\subset X$ is a smooth spacelike Cauchy surface of $X$.
1.9 Proposition.
If $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold, then the associated spacetime $(X,g)$ is a strongly refocusing spacetime with $p=(x,0)$ and $q=(y,l)$. If $(M,h)$ is a $Z^{(x,y)}$ manifold, then the associated spacetime $(X,g)$ is an observer-refocusing spacetime with $p=(x,0)$ and $\gamma:\mathbb{R}\to X,$ $\gamma(t)=(y,t)$.
Proof.
The results follow immediately after realizing that, up to orientation-preserving, affine reparameterization, any future-oriented null-geodesics $\alpha$ in $(X,g)$ through $(x,0)$ is given by $\alpha(t)=(\beta(t),t)$ where $\beta$ is a unit-speed geodesic through $x$ in $(M,h)$. ∎
By proving theorems about the topology of Cauchy surfaces of globally hyperbolic observer-refocusing (resp. strongly refocusing) spacetimes, we will recover theorems about the topology of $Z^{(x,y)}$ (resp. $Y^{(x,y)}_{l}$) manifolds as a special case.
2 Main Results
We briefly state the main results of this paper.
2.10 Result (Theorem 4.37).
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$, let $p\in X$, let $L$ be some lightsphere at $p$ (see Definition 4.26), let $\{\gamma_{i}\}_{i\in I}$ be some countable family of timelike curves and let $T>0$. If for each future-oriented lightlike geodesic $\alpha$ with $\alpha(0)=p$ and $\alpha^{\prime}(0)\in L$ there is some $t\in(0,T]$ and $i\in I$ such that $\alpha(t)\in\mathrm{im}(\gamma_{i})$, then $(X,g)$ has compact Cauchy surface and finite fundamental group.
2.11 Result (Proposition 4.38).
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$ which is observer-refocusing with respect to a point $p\in X$ and a timelike curve $\gamma$ such that the domain of $\gamma$ is compact, then $(X,g)$ has compact Cauchy surface and finite fundamental group.
2.12 Result (Corollary 4.39).
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ such that all unit-speed geodesics starting from $x$ reach $y$ in uniformly bounded time. Then $M$ is compact with finite fundamental group.
2.13 Result (Theorem 5.44).
Let $(X,g)$ be a globally hyperbolic, observer-refocusing spacetime of dimension at least $3$, with $g$ an analytic metric. Then $(X,g)$ is strongly refocusing.
2.14 Result (Corollary 5.45).
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ with $h$ an analytic metric, then $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold for some $l>0$.
3 Related results about $Y^{x}_{l}$ manifolds and weaker notions of refocusing spacetimes
We first review results about the topology of $Y^{x}_{l}$ manifolds and refocusing spacetimes. $Y^{x}_{l}$ manifolds, as well as $Z^{x}$ manifolds and a whole family of related notions, are discussed in [8, Definitions 7.7]. The following result is due to Bott [9] and Samelson [26].
3.15 Theorem (Bott-Samelson).
Let $(M,h)$ be a complete Riemannian manifold of dimension at least $2$, $x\in M$ and $l>0$ such that all unit-speed geodesics $\gamma$ with $\gamma(0)=x$ are periodic with period $l$. Then $\pi_{1}(M)$ is finite and the integral cohomology ring of $M$ is generated by one element.
In [2] Bérard-Bergery proves a similar statement only requiring the geodesics to be closed (as opposed to periodic).
3.16 Theorem (Bérard-Bergery).
Let $(M,h)$ be a $Y^{x}_{l}$ manifold of dimension at least $2$, then $M$ is compact, has finite fundamental group and the rational cohomology ring of $M$ is generated by one element.
By recent contact Bott-Samelson Theorem results of Frauenfelder, Labrousse and Schlenk [13], the conclusion of the Bérard-Bergery theorem can be strengthened to be exactly the same as the one in Bott-Samelson Theorem. Note that is it not know if every $Y^{x}_{l}$ manifold admits a metric in which all geodesics at $x$ are periodic.
On the Lorentzian side, in [19], Low defines a weaker version of a strongly refocusing spacetime.
3.17 Definition (refocusing spacetime).
Let $(X,g)$ be a Lorentzian manifold. We say that $(X,g)$ is refocusing if there exists some $p\in X$ and some open set $U\ni x$ such that for all $V$ with $x\in V\subset U$ there is some $y\in X\setminus V$ such that all null-geodesics through $y$ enter $V$.
Low proved that globally hyperbolic spacetimes must have compact Cauchy surface. In [12, Theorem 11.5] Chernov and Rudyak prove that globally hyperbolic refocusing spacetimes of dimension at least $3$ have a globally hyperbolic semi-Riemannian universal cover which is also refocusing. They use this to prove the following result.
3.18 Theorem.
Every refocusing, globally hyperbolic spacetime of dimension at least $3$ has compact Cauchy surface and finite fundamental group.
In [10] Chernov, Kinlaw and Sadykov define a Riemannian analogue of refocusing spacetimes.
3.19 Definition ($\tilde{Y}^{x}$ manifolds).
Let $(M,h)$ be a complete Riemannian manifold and let $x\in M$. We say that $M$ is a $\tilde{Y}^{x}_{l}$ manifold, if there exists some $\bar{\varepsilon}>0$ such that for all $\varepsilon$ with $0<\varepsilon<\bar{\varepsilon}$ there is some $l>\varepsilon$ and some $y$ such that for all unit-speed geodesics $\gamma$ starting at $y$ we have $d(x,\gamma(l))<\varepsilon$.
3.20 Proposition.
If $(M,h)$ is a $\tilde{Y}^{x}$ manifold, then the associated spacetime $(X,g)$ is refocusing.
The authors combine these results to show that every $\tilde{Y}^{x}$ manifold of dimension at least $2$ is compact with finite fundamental group.
3.21 Remark.
Just as it is unknown whether there are $Z^{x}$ manifolds that are not $Y^{x}_{l}$ manifolds [8, Question 7.70], it is unclear whether there are $\tilde{Y}^{x}$ manifolds that are not $Y^{x}_{l}$ manifolds, whether there are refocusing spacetimes that are not strongly refocusing, and whether there are observer-refocusing spacetimes that are not strongly refocusing.
It is tempting to define some notion of $\tilde{Z}^{x}$ manifold and a corresponding notion for spacetimes.
3.22 Definition ($\tilde{Z}^{x}$ manifolds).
Let $(M,h)$ be a complete Riemannian manifold, let $x\in M$. We say that $(M,h)$ is a $\tilde{Z}^{x}$ manifold if there exists some $\overline{\varepsilon}>0$ such that for all $\varepsilon$ with $0<\varepsilon<\overline{\varepsilon}$ and all unit-speed geodesics $\gamma$ starting at $x$ there is some $t>\varepsilon$ with $d(x,\gamma(t))<\varepsilon$.
3.23 Definition (very weakly refocusing spacetime).
Let $(X,g)$ be a spacetime and $p\in X$. We say that $(X,g)$ is very weakly refocusing if every future-pointing null-geodesic starting at $x$ eventually enters the timelike future $I^{+}(p)$.
Every $Z^{x}$ manifold is a $\tilde{Z}^{x}$ manifold but the converse is false. For example, any $n$-torus $T^{n}$ with standard metric will be a $\tilde{Z}^{x}$ manifold, but only $T^{1}$ is a $Z^{x}$ manifold. Similarly there are not observer-refocusing. We will show in Proposition 4.29 that globally hyperbolic very weakly refocusing spacetimes have compact Cauchy surface. In fact every proof of compactness of Cauchy surfaces in this paper will ultimately rely on this face.
3.24 Proposition.
Let $(M,h)$ be a $\tilde{Z}^{x}$ manifold. Then the associated spacetime $(X,g)$ is very weakly refocusing.
Proof.
Let $\alpha$ be a future-pointing null-geodesic with $\alpha(0)=(x,0)\eqqcolon p$. We can parametrize $\alpha$ so that $\alpha(t)=(\beta(t),t)$ for some unit-speed geodesic $\beta$ in $M$ with $\beta(0)=x$. Since $(M,h)$ is a $\tilde{Z}^{x}$ manifold, there is some $\varepsilon>0$ and some $T>\varepsilon$ such that $d(x,\beta(t))<\varepsilon$. We can assume that $d(x,\beta(T))\neq 0$ else increase $T$ by a tiny bit. Let $\tilde{\beta}$ be a unit-speed geodesic with $\tilde{\beta}(0)=x$ and $\tilde{\beta}(d(x,\beta(T))=\beta(T)$. Consider the curve $\gamma$ in $X$ defined via $\gamma(t)=(\tilde{\beta}(t),\frac{\varepsilon}{d(x,\beta(T))}t)$. It is easy to see that $\gamma$ is a timelike curve with $\gamma(0)=p$ and $\gamma(d(x,\beta(T))=(\beta(T),\varepsilon)$. Since $T>\varepsilon$ we have that $\alpha(T)=(\beta(T),T)$ is in the timelike future of $(\beta(T),\varepsilon)$, which is in the timelike future of $p=\alpha(0)=\gamma(0)$. This shows that $\alpha$ eventually enters the $I^{+}(p)$. Since this is true for all future-pointing null-geodesics $\alpha$ starting at $p$, we conclude that $(X,g)$ is very weakly refocusing. ∎
3.25 Remark.
We will prove in Corollary 4.28 that every $\tilde{Z}^{x}$ manifold is compact. Further, taking any compact Riemannian manifold $(M,h)$ and any $x\in M$, the Poincaré recurrence theorem guarantees that almost all unit-speed geodesics starting at $x$ will exhibit the recurring behavior described in Definition 3.22. It seems likely that every compact manifold $M$ can be equipped with some Riemannian metric such that there is some $x\in M$ making it into a $\tilde{Z}^{x}$ manifold. Maybe this is even true for a generic Riemannian metric on $M$.
4 Observer-Refocusing Spacetimes
In a Riemannian manifold, normalizing non-zero tangent vectors gives a canonical representative for each spacelike direction at a point. There is no canonical choice of representative for future-pointing null-directions at a point in a Lorentzian manifold, but it will often be useful to make some (smooth) choice of representative for each null-direction.
4.26 Definition (lightsphere).
Let $X$ be a Lorentzian manifold and $p\in X$. We denote by $\mathcal{L}_{p}$ the set of future-pointing null-directions at $p$, i.e. we define $\mathcal{L}_{p}=\{[v]_{\sim}\mid v\text{ future pointing and lightlike}\}$ where $\sim$ is the equivalence relation on the set of null-vectors defined through $v_{1}\sim v_{2}\iff\exists\lambda>0:v_{1}=\lambda v_{2}$. We call a smooth submanifold $L\subset T_{p}X$ a lightsphere at $p$ if every $v\in L$ is lightlike and if the natural map from $L$ to $\mathcal{L}_{p}$ is a homeomorphism.
Such lightspheres always exist, for example one can take any auxiliary Riemannian metric $h$ on a Lorentzian manifold $(X,g)$ and define $L=\{v\in T_{p}X\mid v\text{ future pointing and lightlike and }h(v,v)=1\}$.
4.27 Lemma.
Let $(X.g)$ be a globally hyperbolic spacetime and very weakly refocusing with respect to some $p\in X$. Then the Cauchy surface of $(X,g)$ is compact.
Proof.
For each $v\in L$ let $t_{v}$ be such that $\alpha_{v}(t_{v})\in I^{+}(p)$. Take some open neighborhood $U_{v}$ around $v$ such that for all $w\in\overline{U_{v}}$ we have that $\alpha_{w}(t_{v})\in I^{+}(p)$. Since $L$ is compact, we can cover it with a finite number of such neighborhoods $U_{v_{1}},\dots,U_{v_{n}}$. Consider the set $K=\bigcup_{i=1}^{n}\{\alpha_{w}(t_{v_{i}})\mid w\in\overline{U_{v_{i}}}\}$, which is compact as a finite union of compact sets. Let $S\subset X$ be a Cauchy surface with $K\subset I^{-}(S)$ and consider the set $J^{+}(p)\cap S$. Obviously $J^{+}(p)\cap S$ is a non-empty, closed subset of $S$. Further, we get that
$\partial_{S}\left(J^{+}(p)\cap S\right)\subseteq\partial_{X}(J^{+}(p))\cap S=\left(J^{+}(p)\setminus I^{+}(p)\right)\cap S=\emptyset.$
where the last equality holds for the following reason. Let $q\in J^{+}(p)\setminus I^{+}(p)$. Then $q$ must lie on a future-oriented null-geodesic through $p$, i.e. $q=\alpha_{v}(t)$ for some $v\in L$ and some $t>0$. Let $i\in\{1,\dots,n\}$ be such that $v\in U_{v_{i}}$. Since we chose $q\notin I^{+}(p)$, we must have $t\leq t_{v_{i}}$. But this means that $q\in I^{-}(S)$, hence $q\notin S$. This shows that $J^{+}(p)\cap S$ has empty boundary in $S$, which means $J^{+}(p)\cap S=S$. Since $J^{+}(p)\cap S$ is always compact, we are done. ∎
4.28 Corollary.
If $(M,h)$ is a $\tilde{Z}^{x}$-manifold, then $M$ is compact. In particular every $Z^{x}$ manifold is compact.
Proof.
This follows from Lemma 4.27 by the fact that the spacetime associated to $(M,h)$ is very weakly refocusing. ∎
4.29 Proposition.
Let $(X,g)$ be a globally hyperbolic, $p\in X$ and $\gamma$ a timelike curve such that $(X,g)$ is observer-refocusing with respect to $p$ and $\gamma$. If $p=\gamma(0)$, then $(X,g)$ is very weakly refocusing. In particular $(X,g)$ has compact Cauchy surface.
Proof.
For every future-directed timelike vector $v\in T_{p}X$ there is some $q\in\mathrm{im}(\gamma)$ and $t>0$ such that $\alpha_{v}(t)\in\mathrm{im}(\gamma)$, i.e. $\alpha_{v}(t)=\gamma(s)$ for some $s\in\mathbb{R}$. We have $\alpha_{v}(t)\in J^{+}(p)\setminus\{p\}$. Since $\gamma$ is timelike we must have $s>0$ and hence $\alpha_{v}(t)=\gamma(s)\in I^{+}(p)$. We apply Lemma 4.27 and are done. ∎
4.30 Lemma.
Let $(X,g)$ be a globally hyperbolic spacetime, $p\in X$, let $L$ be a lightsphere at $p$, let $A\subseteq L$ be a dense subset of $L$, and let $T>0$. If for every null-geodesic $\alpha$ with $\alpha(0)=p$ and $\alpha^{\prime}(0)\in A$ we have that $\alpha(T)\in I^{+}(p)$, then $X$ has compact Cauchy surface.
Proof.
We can assume that $(X,g)$ is null-geodesically complete, else consider a globally hyperbolic spacetime conformally equivalent to $(X,g)$ that is. Let $v\in L$ and let $v_{n}$ be a sequence in $A$ converging towards $v$ Let $t_{n}$ be such that for every $n\in\mathbb{N}$ we have that $\alpha_{v_{n}}(t_{n})$ is the first null-cut point of $\alpha|_{[0,t_{n}]}$. By assumption the sequence $(t_{n})_{n\in\mathbb{N}}$ is bounded. Without loss of generality assume that the sequence converges, else go to a subsequence. Since the null-cut locus $C_{N}^{+}(p)$ is closed (see [1]), we have $\alpha_{v}(t)\in C_{N}^{+}(p)$. Therefore $\alpha_{v}(t^{\prime})\in I^{+}(p)$ for all $t^{\prime}>t$. The result follows by applying Lemma 4.27. ∎
4.31 Lemma.
Let $M$ be a smooth manifold and let $(x_{n})_{n\in\mathbb{N}}$ be some sequence with some point $x\in M$ as a limit point. Then there exists a $\mathcal{C}^{1}$ curve $\gamma:[0,a]\to M$ with $\gamma(0)=x$, $\gamma^{\prime}(0)\neq 0$ and such that for every $\varepsilon>0$ there are infinitely many distinct points $x_{n}$ in $\gamma([0,\varepsilon])$.
Proof.
Since the result is entirely local, we can assume $M=\mathbb{R}^{k}$ and $x=0$. Without loss of generality, assume that $x_{n}\to x$ and that $x_{n}\neq x$ for all $n$, else take a subsequence. Further, without loss of generality, assume that $\lim_{n\to\infty}\frac{x_{n}}{||x_{n}||}=v$ for some unit vector $v$ and that $||x_{n+1}||<\frac{1}{2^{n}}||x_{n}||$ for all $n$, else we again take a subsequence. Then the sequence $(w_{n})_{n\in\mathbb{N}}$ defined through $w_{n}\coloneqq\frac{x_{n}-x_{n+1}}{||x_{n}-x_{n+1}||}$ converges to $v$. We define the curve $\gamma$ as follows. We connect $x_{n+1}$ to $x_{n}$ by a a $\mathcal{C}^{1}$ curve $\gamma_{n}:[0,a_{n}]$, parametrized by arc-length, satisfying $\gamma_{n}(0)=x_{n+1}$, $\gamma_{n}(a_{n})=x_{n}$, $\gamma_{n}^{\prime}(0)=w_{n+1}$ and $\gamma_{n}^{\prime}(a_{n})=w_{n}$. Since the sequence $(w_{n})_{n\in\mathbb{N}}$ converges to $v$, the interpolations can be done in a way (for example by interpolating with cubic splines) such that that $\lim_{n\to\infty}\sup_{t\in[0.a_{n}]}||\gamma_{n}^{\prime}(t)-v||=0$ and that $\sum_{n\in\mathbb{N}}a_{n}<\infty$. We then define $\gamma$ as the infinite concatenation
$\gamma:\left(0,\sum_{n\in\mathbb{N}}a_{n}\right]\to\mathbb{R}^{k},$
$\gamma\coloneqq\dots\gamma_{n+1}\cdot\gamma_{n}\cdot\gamma_{n-1}\dots\gamma_{1}\cdot\gamma_{0}$
By construction $\gamma$ is $\mathcal{C}^{1}$, as we always have $\gamma_{n+1}^{\prime}(a_{n+1})=w_{n+1}=\gamma_{n}(0)$. Further we have $\lim_{t\to 0}\gamma(t)=x$ as well as $\lim_{t\to 0}\gamma^{\prime}(t)=v$, so $\gamma$ extends to a $\mathcal{C}^{1}$ curve defined on $[0,\sum_{n\in\mathbb{N}}a_{n}]$ with the desired property.
4.32 Lemma.
Let $(X,g)$ be a Lorentzian manifold. $h:[0,1]\to T_{p}X$ a curve such that $h(t)$ is lightlike for each $t\in[0,1]$. Then $h^{\prime}(0)$ is not timelike. Hence $d\exp_{0}(h^{\prime}(0))=\frac{d}{dt}|_{t=0}\exp(h(t))$ is not timelike.
Proof.
We have that $g(h(t),h(t))=0$ for all $t\in[0,1]$. Therefore $02g(h(0),h^{\prime}(0))=\frac{d}{dt}|_{t=0}g(h(t),h(t))=$ 0. This means that $h^{\prime}(0)$ is perpendicular to the null-vector $h(0)$, and hence $h^{\prime}(0)$ cannot be timelike. ∎
4.33 Lemma.
Let $X$ be a globally hyperbolic spacetime, $p\in X$, and $L$ an embedded lightsphere at $p$. If $(v_{n},t_{n})_{n\in\mathbb{N}}$ is a sequence in $L\times\mathbb{R}$ such that $\alpha_{v_{n}}(t_{n})$ is always defined and such that $(\alpha_{v_{n}}(t_{n}))_{n\in\mathbb{N}}$ converges to some $q\in X$, then $(t_{n})_{n\in\mathbb{N}}$ is bounded.
Proof.
Let $\mathcal{T}:X\to\mathbb{R}$ be a smooth, surjective temporal function such that every level set of $\mathcal{T}$ is a Cauchy surface. Take some $s_{1},s_{2}\in\mathbb{R}$ such that $s_{1}<\mathcal{T}(q)<s_{2}$. Since $X$ is globally hyperbolic and since level sets of $\mathcal{T}$ are Cauchy surfaces, for every $v\in L$ and $s\in[s_{1},s_{2}]$, the equation $\mathcal{T}(\exp(tv))=s$ is satisfied by a unique $t\in\mathbb{R}$. Since $\mathcal{T}$ has timelike gradient, we have $\frac{d}{dt}\mathcal{T}(\exp tv)\neq 0$ for all $t$. By the implicit function theorem, the function $f:L\times[s_{1},s_{2}]\to\mathbb{R}$ implicitly defined through $\mathcal{T}(\exp(f(v)v))=s$ is smooth. Since $L\times[s_{1},s_{2}]$ is compact, so is its image under $f$. Let $N\in\mathbb{N}$ be such that for all $n\geq N$ we have $\mathcal{T}(\alpha_{v_{n}}(t_{n}))\in(s_{1},s_{2})$. This means that for all $n\geq N$ we have $t_{n}\in f(L\times[s_{1},s_{2}])$, which is compact hence bounded. ∎
4.34 Proposition.
Let $(X,g)$ be a globally hyperbolic spacetime, $p\in X$ and let $\gamma$ be a timelike curve in $X$. Then the set $R_{p}^{\gamma}\coloneqq\{q\in\mathrm{im}(\gamma)\mid\exists v\in T_{p}X\text{ lightlike and future pointing and }t>0\text{ with }\alpha_{v}(t)=q\}$ is discrete.
Proof.
We can assume that $(X,g)$ is null-geodesically complete, else consider some conformally equivalent spacetime that is. Let $L$ be an embedded lightsphere at $p$. Assume that there is a sequence $(q_{n})_{n\in\mathbb{N}}$ with $q_{n}\in R_{p}^{\gamma}$ converging to some $q\in X$ with $q_{n}\neq q$ for all $n\in\mathbb{N}$. For each $n\in\mathbb{N}$, let $(v_{n},t_{n})\in L\times\mathbb{R}$ be such that $\alpha_{v_{n}}(t_{n})=q_{n}$. By Lemma 4.33, we know that $(t_{n})_{n\in\mathbb{N}}$ is bounded. Without loss of generality, assume that $(v_{n},t_{n})_{n\in\mathbb{N}}$ converges to some $(v,t)\in L\times\mathbb{R}$, else use compactness of $L$ and boundedness of $(t_{n})_{n\in\mathbb{N}}$ to go to a convergent subsequence.
Let $\beta:[0,1]\to L\times\mathbb{R}$ be a $\mathcal{C}^{1}$ curve with $\beta(0)=(v,t)$ and $\beta^{\prime}(0)\neq 0$ and such that for every $\varepsilon>0$, there are infinitely many $(v_{n},t_{n})\in\beta([0,1])$. For $s\in[0,1]$, let us write $\beta(s)=(v_{s},t_{s})$, abusing the notation slightly by giving new meaning to the terms to $v_{0},v_{1}$ and $t_{0},t_{1}$.
Consider the curve $\alpha(s)$ defined through $\alpha(s)\coloneqq\exp(t_{s}v_{s})=\alpha_{v_{s}}(t_{s})$. $\alpha$ is clearly $\mathcal{C}^{1}$. Further, for every $\varepsilon>0$, there are infinitely many intersections of $\alpha|_{[0,\varepsilon]}$ and $\gamma$. Let $t$ be such that $\alpha(0)=q=\gamma(t)$. Since both $\alpha$ and $\gamma$ are $\mathcal{C}^{1}$, the fact that $\alpha|_{[0,\varepsilon]}$ intersects $\gamma$ an infinite number of times for every $\varepsilon>0$ implies that $\alpha^{\prime}(0)$ and $\gamma^{\prime}(t)$ are linearly dependent. Since by construction $\alpha^{\prime}(0)\neq 0$, there is some $\lambda\neq 0$ with $\alpha^{\prime}(0)=\lambda\cdot\gamma^{\prime}(t)$. This is a contradiction to the fact that $\alpha^{\prime}(0)$ is not timelike due to Lemma 4.32. ∎
4.35 Corollary.
Let $(M,h)$ be a Riemannian manifold, $x,y\in M$. Then $R_{x}^{y}\coloneqq\{t>0\mid\text{ there exists a unit-speed geodesic }\gamma\text{ with }\gamma(0)=x\text{ and }\gamma(t)=y\}$ is a discrete set.
Proof.
Consider the associated spacetime $(X,g)$ of $(M,h)$ and set $p=(x,0)$ and $\gamma(t)=(y,t)$. Then $R_{x}^{y}=\gamma^{-1}(R_{p}^{\gamma})$ is discrete since $R_{p}^{\gamma}$ is by Proposition 4.34. ∎
4.36 Proposition.
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$, let $p\in X$, let $L$ be some lightsphere at $p$, let $\{q_{i}\}_{i\in I}\subset X$ be some countable set of points in $X$ and let $T>0$. If for each future-oriented lightlike geodesic $\alpha$ with $\alpha(0)=p$ and $\alpha^{\prime}(0)\in L$ there is some $t\in(0,T]$ and $i\in I$ such that $\alpha(t)\in q_{i}$, then the Cauchy surface of $(X,g)$ is compact and has finite fundamental group.
Proof.
We can assume that $(X,g)$ is null-geodesically complete, else consider a conformally equivalent spacetime that is. For each $i\in I$, consider $K_{i}\coloneqq\{v\in L\mid\exists t\in(0,T]\mid\alpha_{v}(t)=q_{i}\}$. Let $v\in\mathrm{int}_{L}(K_{i})$. Then there is a (non-constant) null-geodesic variation through $v$, with each null-geodesic going through $p$ and $q_{i}$. This means that $q_{i}$ is conjugate to $p$ along $\alpha_{v}|_{[0,T]}$. Hence $\alpha_{v}(t)\in I^{+}(p)$ for all $t>T$. Let $A=L\setminus\bigcup_{i\in I}\partial_{L}(K_{i})$, then $A$ is dense by the Baire category theorem. Further, if $v\in A$, then there is some $i\in I$ with $v\in K_{i}\setminus\partial_{L}(K_{i})=\mathrm{int}_{L}(K_{i})$, and hence $\alpha_{v}(t)\in I^{+}(p)$ for all $t\geq T$. Since $A$ is dense, $X$ has compact Cauchy surface.
To show that the Cauchy surface of $(X,g)$ has finite fundamental group, we proceed as follows. Let $\rho:(\tilde{X},\tilde{g})\to(X,g)$ be the semi-Riemannian universal cover of $X$. Let $\tilde{p}$ be any point in $\rho^{-1}(p)$ and let $\tilde{L}$ be the lightsphere at $p$ given by $\tilde{L}=(d\rho)^{-1}(L)$. Note that the set $\rho^{-1}(\{q_{i}\}_{i\in I}\})$ is countable. Let $\tilde{v}\in\tilde{L}$ and $v=d\rho(\tilde{v})\in L$. Then there is some $t\in(0,T]$ and $i\in I$ with $\alpha_{v}(t)=q_{i}$. Lifting $\alpha_{v}$ to $\tilde{X}$, we get that $\alpha_{\tilde{v}}(t)\in\rho^{-1}(q_{i})\in\rho^{-1}(\{q_{i}\}_{i\in I}\})$. Note that $\rho^{-1}(\{q_{i}\}_{i\in I}\})$ is countable. Through application of the first part of this theorem, we can conclude that $\tilde{X}$ has compact Cauchy surface as well. Now pick a Cauchy surface $S\subset X$ and let $\tilde{S}=\rho^{-1}(S)$. Then $\tilde{S}$ is a Cauchy surface of $\tilde{X}$ (see the proof of [12, Theorem 11.5]) and hence compact. Further, $\tilde{S}$ is simply connected (as $\tilde{X}$ is simply connected and diffeomorphic to $\tilde{S}\times\mathbb{R}$), and $\rho:\tilde{S}\to S$ is a covering map. Pick any $x\in\tilde{S}$, then
#$\#\pi_{1}(S,x)=\#\rho^{-1}(x),$
which is finite since $\tilde{S}$ is compact. ∎
4.37 Theorem.
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$, let $p\in X$, let $L$ be some lightsphere at $p$, let $\{\gamma_{i}\}_{i\in I}$ be some countable family of timelike curves and let $T>0$. If for each future-oriented lightlike geodesic $\alpha$ with $\alpha(0)=p$ and $\alpha^{\prime}(0)\in L$ there is some $t\in(0,T]$ and $i\in I$ such that $\alpha(t)\in\mathrm{im}(\gamma_{i})$, then $(X,g)$ has compact Cauchy surface and finite fundamental group.
Proof.
For each $i\in I$ the set $R_{p}^{\gamma_{i}}$ is discrete, hence countable. The claim follows immediately from applying the Proposition 4.36 to the countable set $\bigcup_{i\in I}R_{p}^{\gamma_{i}}$. ∎
4.38 Proposition.
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$ which is observer-refocusing with respect to a point $p\in X$ and a timelike curve $\gamma$ such that the domain of $\gamma$ is compact, then $(X,g)$ has compact Cauchy surface and finite fundamental group.
Proof.
Using similar arguments as in Lemma 4.33 compactness of $\mathrm{im}(\gamma)$ ensures that there is some lightsphere $L$ at $p$ and some $T>0$ such that for all $v\in L$ there is some $t\in(0,T]$ with $\gamma(t)\in\mathrm{im}(\gamma)$. Now we can use Theorem 4.37
4.39 Corollary.
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ such that all unit-speed geodesics starting from $x$ reach $y$ in uniformly bounded time. Then $M$ is compact with finite fundamental group.
Proof.
Let $T>0$ be the uniform bound on the time it takes unit-speed geodesics starting at $x$ to reach $y$. Consider the associated spacetime $(X,g)$ to $(M,h)$. Defining $p=(x,0)$ and $\gamma:[0,T]\to X$, $\gamma(t)=(y,t)$, the result then follows from Proposition 4.38. ∎
4.40 Corollary.
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension $2$ or $3$, let $x\in M$ and assume that all unit-speed geodesic starting at $x$ reach $y$ in uniformly bounded time. Then there exists a Riemannian metric $\tilde{h}$ on $M$ such that $(M,\tilde{h})$ is a $Y^{x}_{l}$ manifold.
Proof.
In this case $M$ is compact and has finite fundamental group. If $\dim(M)=3$, then by the Thurston Elliptization conjecture [27] proved by Perelman [23], [24], [25], $M$ is a quotient of the sphere $S^{3}$ (with standard metric) by a finite group of isometries. The induced metric on $M$ will make it into a $Y^{x}_{2\pi}$ manifold. Similar arguments work for $\dim(M)=2$. ∎
4.41 Remark.
If we drop the requirement of a uniform bound from Theorem 4.37 we can still show that there is an open, dense subset $A\subset L$ such that all null-geodesics $\alpha$ starting at $p$ with initial velocity in $A$ eventually enter $I^{+}(p)$, but this is not enough to conclude that we have a compact Cauchy surface.
4.42 Remark.
It is unknown whether there are $Z^{x}$ manifolds where the time-of-return of unit-speed geodesics is not uniformly bounded [8, Question 7.71]. Similarly it is unknown whether there are any globally hyperbolic observer-refocusing spacetimes where the timelike curve cannot be taken to be compact.
5 The Analytic Case
If our metric is analytic instead of just smooth, an observer-refocusing spacetime of dimension at least $3$ is always strongly refocusing.
5.43 Lemma.
Let $(X,g)$ be a globally hyperbolic spacetime with $g$ an analytic metric. Then there exists an analytic, surjective temporal function $\mathcal{T}:X\to\mathbb{R}$ such that every level set of $\mathcal{T}$ is a spacelike Cauchy hypersurface.
Proof.
Since $X$ is globally hyperbolic, Bernal and Sanchez proved in [3], [4] that there exists an isometry between $(X,g)$ and $(S\times\mathbb{R},\hat{g})$, where $S$ is a Cauchy surface of $X$ and $\hat{g}$ is a metric of the form $\hat{g}=h-Adt^{2}$ for some smooth function $A:S\times\mathbb{R}\to\mathbb{R}^{+}$ and so that $h$ is a symmetric tensor which restricts to a Riemannian metric on every hypersurface $S\times\{t\}$ and with $h(\partial_{t},\cdot)=0$. Let $\hat{\mathcal{T}}:S\times\mathbb{R}\to\mathbb{R}$ be the projection onto the time-coordinate, then $\hat{\mathcal{T}}$ is an analytic, surjective temporal function with all level sets spacelike Cauchy surfaces. By a theorem of Myers and Steenrod [20], isometries are always at least as regular as the metric. Therefore the isometry $\varphi:(X,g)\to(S\times\mathbb{R},\hat{g})$ is analytic. The temporal function $\mathcal{T}\coloneqq\hat{\mathcal{T}}\circ\phi:X\to\mathbb{R}$ is then analytic as the composition of two analytic functions. ∎
5.44 Theorem.
Let $(X,g)$ be a globally hyperbolic, observer-refocusing spacetime of dimension at least $3$, with $g$ an analytic metric. Then $(X,g)$ is strongly refocusing.
Proof.
Let $p\in X$ and let $\gamma$ be a timelike curve in $X$ so that $(X,g)$ is observer-refocusing with respect to $p$ and $\gamma$. Let $L\subset T_{p}X$ be some lightsphere at $p$ which is also an analytic submanifold of $T_{p}X$, i.e. a lightsphere such that the inclusion $\iota:L\to T_{p}X$ is analytic. Let $D\subset L\times\mathbb{R}$ be the maximal domain of the function $(t,v)\mapsto\exp_{p}(tv)$. Note that the map $H:D\to T_{p}X$ defined through $H(v,t)=t\cdot\iota(v)$ is analytic. Let $\mathcal{T}$ be a surjective, analytic temporal function such that each level set of $\mathcal{T}$ is a Cauchy surface and consider the map
$\phi:D\to\mathbb{R}$
$\phi(v,t)=\mathcal{T}(\exp_{p}(tv))$
Since $g$ is analytic, so is the exponential map $\exp_{p}:T_{p}X\to X$. Hence $\phi=\mathcal{T}\circ\exp_{p}\circ H$ is analytic as the composition of analytic maps.
For each $q\in R_{p}^{\gamma}$ consider the set $K_{q}\coloneqq\{v\in L\mid\exists t>0\text{ with }\alpha_{v}(t)=q\}$. Each $K_{q}$ is a closed set and $L=\bigcup_{q\in R_{p}^{\gamma}}K_{q}$ is a countable union of these sets. By the Baire category theorem, there has to be some $q\in R_{p}^{\gamma}$ such that $U\coloneqq\mathrm{int}_{L}(K_{q})$ is non-empty. Let $T=\mathcal{T}(q)$. Then, for every $v\in L$, there is a unique $t\in\mathbb{R}$ with $(v,t)\in D$ and $\phi(v,t)=T.$ By the analytic implicit function theorem, there is an analytic map $f:L\to\mathbb{R}$ with $(v,f(v))\in D$ and $\phi(v,f(v))=T$ for all $v\in L$. Now consider the map
$\psi:L\to X$
$\psi(v)=\exp_{p}(H(f(v),v))=\exp_{p}(f(v)\cdot v)$
which is analytic since the maps $\exp_{p}$, $H$, and $f$ are analytic. For $v\in U$ we must have $\phi(v)=q$, so $\psi$ is constant on $U$. Since $\psi$ is analytic on all of $L$, $\psi$ must be constant on $L$. ∎
5.45 Corollary.
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ with $h$ an analytic metric, then $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold for some $l>0$.
In particular, any $Z^{x}$ manifold with analytic metric is a $Y^{x}_{l}$ manifold for some $l>0$ (the $1$-dimensional case is trivially true even without analyticity).
6 Analogous conjecture in the contact case
The results of Chernov and Nemirovski [11], the results of Frauenfelder, Labrousse and Schlenk [13, Theorem 1.11], and our results about observer-refocusing spacetimes and Riemannian $Z^{x}$ manifolds in this paper, suggest the following conjectures.
6.46 Conjecture.
Let $M$ be manifold of dimension at least $2$, let $(S^{\ast}M,\xi)$ be the spherical cotangent bundle on $M$ equipped with the standard contact structure. Let $\alpha$ be any contact form on $(S^{\ast}M,\xi)$, i.e. $\xi=\ker(\alpha)$, and let $\varphi_{\alpha}^{t}$ be the Reeb flow of $\alpha$. Assume that for all $v\in S^{\ast}_{x}M$ there is some $t>0$ with $\varphi_{\alpha}^{t}(v)\in S^{\ast}_{x}M$. Then $M$ is compact, the fundamental group of $M$ is finite, and the integral cohomology ring of the universal cover of $M$ is the one of a CROSS.
6.47 Conjecture.
Let $M$, $\xi$, $\alpha$ and $\varphi_{\alpha}^{t}$ be as in Conjecture 6.46. Let $x\in M$. Assume there is some $T>0$ such that for all $v\in S^{\ast}_{x}M$ there is some $t\in(0,T]$ with $\varphi_{\alpha}^{t}(v)\in S^{\ast}_{x}M$. Then $M$ is compact, the fundamental group of $M$ is finite and the integral cohomology ring of the universal cover of $M$ is the one of a CROSS.
This contact-geometric version generalizes the Riemannian $Z^{x}$ version as follows: Let $(M,g)$ be a Riemannian manifold, let $\psi:TM\to T^{\ast}M$ be the isomorphism defined through $v\mapsto g(v,\cdot)$, then $\psi$ induces a bundle-isomorphism (which we will also denote by $\psi$) between the unit-tangent bundle $U_{g}M$ and the spherical cotangent bundle $S^{\ast}M$. We can define a $1$-form $\alpha$ on $S^{\ast}M$ via $\alpha_{v}(w)=g(\psi^{-1}(v),d\pi(w))$ where $\pi:S^{\ast}M\to M$ is the natural projection. The Reeb flow of $\alpha$ on $S^{\ast}M$ corresponds to the co-geodesic flow of $g$ on $S^{\ast}M$. Conjecture 6.46 generalizes the conjecture that $Z^{x}$ manifolds of dimension at least $2$ have finite fundamental group. We have proven the compactness and the fundamental group aspect of Conjecture 6.47 whenever the $1$-forms $\alpha$ comes from a Riemannian metric. The conjecture states that this is really a fact about Reeb flows on $(S^{\ast}M,\xi)$ and not just about geodesics on $M$. One may also ask other questions, for example if the set $R_{x}^{\alpha}\coloneqq\{t>0\mid\exists v\in S^{\ast}_{x}M\text{ such that }\varphi_{\alpha}^{t}(x)\in S^{\ast}_{x}M\}$ is discrete for all contact forms $\alpha$ on $(S^{\ast}M,\xi)$. If it is always discrete, then Conjecture 6.46 holds for analytic $1$-forms $\alpha$ by the analyticity of the Reeb flow and by the analyticity of $\pi:S^{\ast}M\to M$ because arguments as in Theorem 5.44 can reduce the problem to the case already solved by Frauenfelder, Labrousse and Schlenk.
Acknowledgments
The author thanks Vladimir Chernov for the fruitful idea of investigating the topology of Lorentzian analogues of $Z^{x}$ manifolds.